OpenStudy (anonymous):

Vocabulary Help. Algebra 2. Medal and Fan! A. Branch B. Continuous Graph C. Discontinuous Graph D. Inverse Variation E. Point of Discontinuity F. Rational Equation G. Rational Expression H. Reciprocal Function I. Removable Discontinuity J. Simplest Form 1. A graph that has no jumps, breaks, or holes. 2. An equation that contains a rational expression. 3. The quotient of two polynomials. 4. A point of discontinuity, a, of function f that you can remove by redefining f at x=a. Doing so fills in a hole in the graph of f with the point (a, f(a)).

OpenStudy (anonymous):

5. A relation represented by the equation of the form xy=k, y=k/x, or x=k/y, where k≠0. 6. A piece of a discontinuous graph. 7. The x-coordinate of a point where the graph of f(x) is not continuous. 8. When the numerator and denominator are polynomials that have no common divisor other than 1. 9. A graph that has jumps, breaks, or holes. 10. Belongs to the family whose parent function is f(x) = 1/x, where x≠0.

OpenStudy (anonymous):

im not sure how to do this sorry

OpenStudy (anonymous):

5. Inverse variation

OpenStudy (anonymous):

1. B 2. F 3.G

OpenStudy (anonymous):

4.I 6. A 7. E 8.J 9 C 10 D

OpenStudy (anonymous):

If 5 is inverse variation, how is 10? @phuong.tong

OpenStudy (anonymous):

Sorry . 10 is H, reciprical function